The circumference of a circle is 80 inches. Find the radius to the nearest tenth of an inch.
Found 2 solutions by rfer, checkley77:
Answer by rfer(3375)

C=3.14d
d/2=r
(80/3.14)/2=12.7 in

Answer by checkley77(9413)

80=2*PI*R
80=2*3.14*R
80=6.28*R
R=80/6.28
R=12.7 INCHES IS THE RADIUS.